A regular square pyramid has an altitude of 15 m and a volume of 125 m3.

1. What is the base edge of a square pyramid?
2. What is the slant height of the pyramid?
3. What is the lateral area in m2?
1
Expert's answer
2020-11-17T17:28:05-0500

We have "h=15\ \text{m}, \ \ V=125 \ \text{m}^3."

1 . Let "a" be the base edge of this square pyramid.

Volume of the pyramid is "V=\frac{1}{3} B h" , where "B=a^2" is area of base.

"V=\frac{1}{3}a^2h"

"a^2=\frac{3V}{h}"

"a=\sqrt{\frac{3V}{h}}= \sqrt{\frac{3\times 125}{15}}=\sqrt{25}=5"

2.Let "l" be the slant height of the pyramid.

Then "l=\sqrt{h^2+\left(\frac{a}{2}\right)^2}=\sqrt{15^2+2.5^2}=\sqrt{925/4}=5\sqrt{37}/2\approx 15.2\ \text{m}"

3.We can find the lateral area using formula "L=\frac{1}{2}Pl" , where "P=4a" is the perimeter of the base, "l" is the slight height.

"L=\frac{4al}{2}=\frac{4\times 5\times 5\sqrt{37}}{2\times 2}=25\sqrt{37}\approx 152\ \text{m}^2" .




Answer: 1. "a=5\ \text{m}" , 2. "l=5\sqrt{37}/{2}\approx 15.2\ \text{m}" , 3. "L=25\sqrt{37}\approx 152\ \text{m}^2" .